# Variational Inference for Sparse Poisson Regression

### v8

@@ Acknowledgments
The authors would like to thank five anonymous reviewers for their valuable comments and corrections, which significantly improved the results of this paper.
@@
$=$ $(\mathbf{y}-M_{\boldsymbol{\xi}})^{\top}\mathbf{X}{\boldsymbol{\beta}}-\frac{1}{2}{\boldsymbol{\beta}}^{\top}\left[S_{X}^{\boldsymbol{\xi}}+\text{diag}({\rm E}({\boldsymbol{\tau}}^{-1}))\right]{\boldsymbol{\beta}}+\text{Const.,}$
@@
$=$ ${(p+\nu-2)}\log{\eta}-{\eta}\left(\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}(\tau_{j})+\delta\right)+\text{Const.},$
@@
$=$ $-\frac{1}{2}\log(\tau_{j})-\frac{1}{2}\dfrac{{\rm E}[\beta_{j}^{2}]}{\tau_{j}}-\frac{{\rm E}[{\eta}]}{2}\tau_{j}+\text{Const.},$
@@
$=$ $-2\log\tau_{0}-\frac{1}{\tau_{0}}\left[\frac{1}{2}{\rm E}(\beta_{0}^{2})+{\rm E}(a^{-1})\right]+\text{Const.},$
@@
$=$ $-2\log a-\dfrac{1}{a}({\rm E}(\tau_{0}^{-1})+A^{-1})+\text{Const.},$
@@
$-\frac{1}{2}{\rm tr}(S_{X}^{\boldsymbol{\xi}}D_{{\boldsymbol{\beta}}}^{L})+\mathbf{y}^{\top}\mathbf{X}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(L)}+\text{Const.},$
where $D_{{\boldsymbol{\beta}}}^{L}={\rm E}({\boldsymbol{\beta}}{\boldsymbol{\beta}}^{\top})={\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(L)}+{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(L)}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(L)}^{\top},$
${\rm E}_{q}(\log p({\boldsymbol{\beta}}|\tau))=$ $-\frac{1}{2}{\rm E}(\log\tau_{0})-\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}(\log\tau_{j})$
$-\frac{1}{2}{\rm E}(\tau_{0}^{-1}){\rm E}(\beta_{0}^{2})-\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}(\tau_{j}^{-1}){\rm E}(\beta_{j}^{2})+\text{Const.},$
${\rm E}_{q}(\log p({\eta}))=$ $(\nu-1){\rm E}(\log{\eta})-\delta{\rm E}({\eta})+\text{Const.}$
$=$ $(\nu-1)\left[\psi({p+\nu-1})-\log\left(\delta+\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}(\tau_{j})\right)\right]$
- $\dfrac{\delta({p+\nu-1})}{\delta+\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}(\tau_{j})}+\text{Const.},$
${\rm E}_{q}(\log p(\tau_{0}|a))=$ ${\rm E}\left(-\frac{1}{2}\log a-\frac{3}{2}\log\tau_{0}-\frac{1}{a\tau_{0}}\right)+\text{Const.}$
$=$ $-\frac{1}{2}\log(A^{-1}+{\rm E}(\tau_{0}^{-1}))-\frac{3}{2}\log(d_{00}^{L}/2+{\rm E}(a^{-1}))$
$-{\rm E}(a^{-1}){\rm E}(\tau_{0}^{-1})+\text{Const.},$
@@
$=$ $-{(p-1)}\log\left(\delta+\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}(\tau_{j})\right)-\frac{1}{2}{\rm E}({\eta}){\sum_{j=1}^{p-1}}{\rm E}(\tau_{j})+\text{Const.},$
@@
$-{\rm E}_{q}(\log q({\eta}))=$ $(p+\nu-1)\log\left(\delta+\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}(\tau_{j})\right)+\text{Const.},$
$-{\rm E}_{q}(\log q(\tau_{0}))=$  $\;2{\rm E}(\log(\tau_{0}))+{\rm E}(\tau_{0}^{-1})\big{(}d_{00}^{L}/2+{\rm E}(a^{-1})\big{)}$
$-\log(d_{00}^{L}/2+{\rm E}(a^{-1}))+\text{Const.}$
$=$ $\log(d_{00}^{L}/2+{\rm E}(a^{-1}))+\text{Const.},$
$-{\rm E}_{q}(\log q(\tau))=$ $\;\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}(\log\tau_{j})+\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}(\tau_{j}^{-1}){d_{jj}^{L}}$
$+\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}({\eta}){\rm E}(\tau_{j})-\frac{1}{4}{\sum_{j=1}^{p-1}}{\rm E}({\eta})/{d_{jj}^{L}}$
$+{\sum_{j=1}^{p-1}}\log K_{1/2}(\sqrt{{\rm E}({\eta}){d_{jj}^{L}}})+\text{Const.},$
@@
$-{\rm E}_{q}(\log q(a))=$  $\;2{\rm E}(\log a)+{\rm E}(a^{-1})\big{(}{\rm E}(\tau_{0}^{-1})+A^{-1}\big{)}+\text{Const.}$
$=$  $\;2\log\big{(}{\rm E}(\tau_{0}^{-1})+A^{-1}\big{)}-\psi(1)+{\rm E}(a^{-1})\big{(}{\rm E}(\tau_{0}^{-1})+A^{-1}\big{)}$
$-\log(A^{-1}+{\rm E}(\tau_{0}^{-1}))+\text{Const.}$
$=$ $\log\big{(}{\rm E}(\tau_{0}^{-1})+A^{-1}\big{)}+\text{Const.}$
@@
$-\dfrac{1}{2c{\tau^{2}}}E\left[\text{tr(diag}(1-Z)){\boldsymbol{\beta}}{\boldsymbol{\beta}}^{T}\right]-\frac{1}{{\tau^{2}}}{\rm E}(a^{-1})$
@@
$=$ $-\dfrac{p+1}{2}\log{\tau^{2}}-\dfrac{1}{{\tau^{2}}}\left[\frac{1}{2}\text{tr(diag}({P^{(C)}})D_{{\boldsymbol{\beta}}}^{C})\right$.
$+\left.\dfrac{1}{2c}\text{tr(diag}(1-{P^{(C)}})D_{{\boldsymbol{\beta}}}^{C})+{\rm E}(a^{-1})\right]+\text{Const.},$
@@
$=$ $2\log a-\dfrac{1}{a}({\rm E}({\tau^{-2}})+A^{-1})+\text{Const.},$
@@
$-\frac{1}{2}{\rm tr}(S_{X}^{\boldsymbol{\xi}}D_{{\boldsymbol{\beta}}}^{C})+\mathbf{y}^{\top}\mathbf{X}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(C)}+\text{Const.},$
@@
$=$ $-\frac{1}{2}{\sum_{j=1}^{p-1}}{P_{j}^{(C)}}\big{(}{\rm E}(\log{\tau^{2}})+{\rm E}({\tau^{-2}}){d_{jj}^{C}}\big{)}$
$-\frac{1}{2}{\sum_{j=1}^{p-1}}(1-{P_{j}^{(C)}})\big{(}{\rm E}(\log{\tau^{2}})+\frac{1}{c}{\rm E}({\tau^{-2}}){d_{jj}^{C}}\big{)}$
$=$ $-\frac{1}{2}(1-\frac{1}{c}){\rm E}({\tau^{-2}}){\sum_{j=1}^{p-1}}{P_{j}^{(C)}}{d_{jj}^{C}}-\frac{1}{2c}{\rm E}({\tau^{-2}}){\sum_{j=1}^{p-1}}{d_{jj}^{C}}$
$-\frac{p-1}{2}{\rm E}(\log{\tau^{2}})+\text{Const.},$
${\rm E}_{q}(\log p({\tau^{2}}|a))=$ $-\frac{1}{2}{\rm E}(\log a)-\frac{3}{2}{\rm E}(\log{\tau^{2}})-\frac{1}{2}{\rm E}({\tau^{-2}}){\rm E}(a^{-1})+\text{Const.},$
@@
$=$ ${\sum_{j=1}^{p-1}}[{P_{j}^{(C)}}E\left(\log{\pi_{j}}\right)+(1-{P_{j}^{(C)}}){\rm E}(\log(1-\pi_{j}))]+\text{Const.},$
${\rm E}_{q}(\log p(\pi))=$ $({\rho}_{1}-1){\sum_{j=1}^{p-1}}{\rm E}(\log\pi_{j})+({\rho}_{2}-1){\sum_{j=1}^{p-1}}{\rm E}(\log(1-\pi_{j}))+\text{Const.},$
@@
$=$ $-\frac{3}{2}{\rm E}(\log a)-A^{-1}{\rm E}(a^{-1})+\text{Const.},$
@@
$+(\alpha_{{\tau^{2}}}-1){\rm E}(\log{\tau^{2}})+\beta_{{\tau^{2}}}E({\tau^{-2}})+\text{Const.},$
@@
$-{\rm E}_{q}(\log q(a))=$ $-\log\Big{(}A^{-1}+{\rm E}({\tau^{-2}})\Big{)}+2{\rm E}(\log a)+\Big{(}A^{-1}+{\rm E}({\tau^{-2}})\Big{)}{\rm E}(a^{-1})+\text{Const.}$
$=$ $\log\Big{(}A^{-1}+{\rm E}({\tau^{-2}})\Big{)}+\text{Const.}$
@@
$=$ $(\mathbf{y}-M_{\boldsymbol{\xi}})^{\top}\mathbf{X}\text{diag}({P^{(B)}}){\boldsymbol{\beta}}-\frac{1}{2}{\boldsymbol{\beta}}^{\top}\left[S_{X}^{\boldsymbol{\xi}}\odot\Omega+{\rm diag}({\rm E}(\alpha))\right]{\boldsymbol{\beta}}+\text{Const.},$
@@
$\Omega={\rm E}({\boldsymbol{\gamma}}{\boldsymbol{\gamma}}^{T})=({P^{(B)}})({P^{(B)}})^{\top}+{\rm diag}({P^{(B)}})({I_{p}}-{\rm diag}({P^{(B)}})).$
@@
$=$ $\gamma_{j}\Big{[}(y-M_{\boldsymbol{\xi}})^{T}X_{j}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(B)j}-\frac{1}{2}S_{X_{jj}}^{\xi}{{d_{jj}^{B}}}$
$-\frac{1}{2}\sum_{i\neq j}P_{i}^{(B)}S_{X_{ij}}^{\boldsymbol{\xi}}d_{ij}^{B}+{\rm E}\big{(}\log\big{(}\dfrac{\pi_{j}}{1-\pi_{j}}\big{)}\big{)}\big{]}+\text{Const.},$
@@
$=$ $-M_{\boldsymbol{\xi}}\mathbf{X}{P^{(B)}}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(B)}-\frac{1}{2}\text{tr}[D_{{\boldsymbol{\beta}}}^{B}(S_{X}^{\boldsymbol{\xi}}\odot\Omega)]$
@@
$=$ $\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}(\log\alpha_{j})-\frac{1}{2}\text{tr}\big{[}D_{{\boldsymbol{\beta}}}^{B}\text{diag}({\rm E}(\alpha))\big{]}\Big{]}+\text{Const.},$
@@
$=$ ${\sum_{j=1}^{p-1}}\big{[}(a_{j}-1){\rm E}(\log\alpha_{j})-b_{j}\frac{a_{j}+1/2}{b_{j}+\frac{d_{jj}}{2}}\big{]}+\text{Const.},$
@@
$=$ ${\sum_{j=1}^{p-1}}\left[{P_{j}^{(B)}}{\rm E}(\log\pi_{j})+(1-{P_{j}^{(B)}}){\rm E}(\log(1-\pi_{j}))\right]+\text{Const.},$
${\rm E}_{q}(\log p(\pi))=$ $({\rho}_{1}-1){\sum_{j=1}^{p-1}}{\rm E}(\log\pi_{j})+({\rho}_{2}-1){\sum_{j=1}^{p-1}}{\rm E}(\log(1-\pi_{j}))+\text{Const.},$
@@
$-{\rm E}_{q}(\log q(\boldsymbol{\alpha}))=$ $-{\sum_{j=1}^{p-1}}(a_{j}-1/2){\rm E}(\log\alpha_{j})-{\sum_{j=1}^{p-1}}(a_{j}+1/2)\log(b_{j}+d_{jj}/2)+\text{Const.}$
@@
Based on the results of the simulation study, we present a guideline for choosing the suitable prior in Table 4. The best model is suggested in each of low- and high-dimensional cases, based on the sparsity, estimation, and prediction evaluation criteria.
For a low-dimensional case, the CS-VB and Laplace-VB methods are the best methods in the sense of sparsity, estimation, and prediction performance. For the high-dimensional cases, the Laplace-VB model performs significantly better than the other VB models, for accurate estimation of the regression coefficients, while both CS-VB and Laplace-VB models perform well for prediction of the response. In this case, the Bernoulli-VB method has the best performance based on the FPR and TPR criterions, while the CS-VB method is also the second best method in this case. In general, Both CS-VB and Laplace-VB methods are the most suggested models in all cases and for all purposes.
@@
Sparsity CS & Laplace Bernoulli & CS
Estimation Laplace & CS Laplace
Prediction Laplace & CS CS & Laplace

### v9

@@ Acknowledgments
The authors would like to thank five anonymous reviewers for their valuable comments and corrections, which significantly improved the results of this paper. The reseach is based upon research funded by Iran National Science Foundation (INSF) under project No. 4048851.
@@
$=$ $(\mathbf{y}-M_{\boldsymbol{\xi}})^{\top}\mathbf{X}{\boldsymbol{\beta}}-\frac{1}{2}{\boldsymbol{\beta}}^{\top}\left[S_{X}^{\boldsymbol{\xi}}+\text{diag}({\rm E}_{q}({\boldsymbol{\tau}}^{-1}))\right]{\boldsymbol{\beta}}+\text{Const.,}$
@@
$=$ ${(p+\nu-2)}\log{\eta}-{\eta}\left(\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}_{q}(\tau_{j})+\delta\right)+\text{Const.},$
@@
$=$ $-\frac{1}{2}\log(\tau_{j})-\frac{1}{2}\dfrac{{\rm E}_{q}[\beta_{j}^{2}]}{\tau_{j}}-\frac{{\rm E}_{q}[{\eta}]}{2}\tau_{j}+\text{Const.},$
@@
$=$ $-2\log\tau_{0}-\frac{1}{\tau_{0}}\left[\frac{1}{2}{\rm E}_{q}(\beta_{0}^{2})+{\rm E}_{q}(a^{-1})\right]+\text{Const.},$
@@
$=$ $-2\log a-\dfrac{1}{a}({\rm E}_{q}(\tau_{0}^{-1})+A^{-1})+\text{Const.},$
@@
$-\frac{1}{2}{\rm tr}(S_{X}^{\boldsymbol{\xi}}D_{{\boldsymbol{\beta}}}^{(L)})+\mathbf{y}^{\top}\mathbf{X}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(L)}+\text{Const.},$
where $D_{{\boldsymbol{\beta}}}^{(L)}={\rm E}_{q}({\boldsymbol{\beta}}{\boldsymbol{\beta}}^{\top})={\boldsymbol{\Sigma}}_{{\boldsymbol{\beta}}(L)}+{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(L)}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(L)}^{\top},$
${\rm E}_{q}(\log p({\boldsymbol{\beta}}|\tau))=$ $-\frac{1}{2}{\rm E}_{q}(\log\tau_{0})-\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}_{q}(\log\tau_{j})$
$-\frac{1}{2}{\rm E}_{q}(\tau_{0}^{-1}){\rm E}_{q}(\beta_{0}^{2})-\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}_{q}(\tau_{j}^{-1}){\rm E}_{q}(\beta_{j}^{2})+\text{Const.},$
${\rm E}_{q}(\log p({\eta}))=$ $(\nu-1){\rm E}_{q}(\log{\eta})-\delta{\rm E}_{q}({\eta})+\text{Const.}$
$=$ $(\nu-1)\left[\psi({p+\nu-1})-\log\left(\delta+\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}_{q}(\tau_{j})\right)\right]$
- $\dfrac{\delta({p+\nu-1})}{\delta+\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}_{q}(\tau_{j})}+\text{Const.},$
${\rm E}_{q}(\log p(\tau_{0}|a))=$ ${\rm E}_{q}\left(-\frac{1}{2}\log a-\frac{3}{2}\log\tau_{0}-\frac{1}{a\tau_{0}}\right)+\text{Const.}$
$=$ $-\frac{1}{2}\log(A^{-1}+{\rm E}_{q}(\tau_{0}^{-1}))-\frac{3}{2}\log(d_{00}^{(L)}/2+{\rm E}_{q}(a^{-1}))$
$-{\rm E}_{q}(a^{-1}){\rm E}_{q}(\tau_{0}^{-1})+\text{Const.},$
@@
$=$ $-{(p-1)}\log\left(\delta+\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}_{q}(\tau_{j})\right)-\frac{1}{2}{\rm E}_{q}({\eta}){\sum_{j=1}^{p-1}}{\rm E}_{q}(\tau_{j})+\text{Const.},$
@@
$-{\rm E}_{q}(\log q({\eta}))=$ $(p+\nu-1)\log\left(\delta+\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}_{q}(\tau_{j})\right)+\text{Const.},$
$-{\rm E}_{q}(\log q(\tau_{0}))=$  $\;2{\rm E}_{q}(\log(\tau_{0}))+{\rm E}_{q}(\tau_{0}^{-1})\big(d_{00}^{L}/2+{\rm E}_{q}(a^{-1})\big)$
$-\log(d_{00}^{L}/2+{\rm E}_{q}(a^{-1}))+\text{Const.}$
$=$ $\log(d_{00}^{L}/2+{\rm E}_{q}(a^{-1}))+\text{Const.},$
$-{\rm E}_{q}(\log q(\tau))=$ $\;\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}_{q}(\log\tau_{j})+\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}_{q}(\tau_{j}^{-1}){d_{jj}^{(L)}}$
$+\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}_{q}({\eta}){\rm E}_{q}(\tau_{j})-\frac{1}{4}{\sum_{j=1}^{p-1}}{\rm E}_{q}({\eta})/{d_{jj}^{(L)}}$
$+{\sum_{j=1}^{p-1}}\log K_{1/2}(\sqrt{{\rm E}_{q}({\eta}){d_{jj}^{(L)}}})+\text{Const.},$
@@
$-{\rm E}_{q}(\log q(a))=$  $\;2{\rm E}_{q}(\log a)+{\rm E}_{q}(a^{-1})\big({\rm E}_{q}(\tau_{0}^{-1})+A^{-1}\big)+\text{Const.}$
$=$  $\;2\log\big({\rm E}_{q}(\tau_{0}^{-1})+A^{-1}\big)-\psi(1)+{\rm E}_{q}(a^{-1})\big({\rm E}_{q}(\tau_{0}^{-1})+A^{-1}\big)$
$-\log(A^{-1}+{\rm E}_{q}(\tau_{0}^{-1}))+\text{Const.}$
$=$ $\log\big({\rm E}_{q}(\tau_{0}^{-1})+A^{-1}\big)+\text{Const.}$
@@
$-\dfrac{1}{2c{\tau^{2}}}E\left[\text{tr(diag}(1-Z)){\boldsymbol{\beta}}{\boldsymbol{\beta}}^{T}\right]-\frac{1}{{\tau^{2}}}{\rm E}_{q}(a^{-1})$
@@
$=$ $-\dfrac{p+1}{2}\log{\tau^{2}}-\dfrac{1}{{\tau^{2}}}\left[\frac{1}{2}\text{tr(diag}({P^{(C)}})D_{{\boldsymbol{\beta}}}^{(C)})\right$.
$+\left.\dfrac{1}{2c}\text{tr(diag}(1-{P^{(C)}})D_{{\boldsymbol{\beta}}}^{(C)})+{\rm E}_{q}(a^{-1})\right]+\text{Const.},$
@@
$=$ $2\log a-\dfrac{1}{a}({\rm E}_{q}({\tau^{-2}})+A^{-1})+\text{Const.},$
@@
$-\frac{1}{2}{\rm tr}(S_{X}^{\boldsymbol{\xi}}D_{{\boldsymbol{\beta}}}^{(C)})+\mathbf{y}^{\top}\mathbf{X}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(C)}+\text{Const.},$
@@
$=$ $-\frac{1}{2}{\sum_{j=1}^{p-1}}{P_{j}^{(C)}}\big({\rm E}_{q}(\log{\tau^{2}})+{\rm E}_{q}({\tau^{-2}}){d_{jj}^{(C)}}\big)$
$-\frac{1}{2}{\sum_{j=1}^{p-1}}(1-{P_{j}^{(C)}})\big({\rm E}_{q}(\log{\tau^{2}})+\frac{1}{c}{\rm E}_{q}({\tau^{-2}}){d_{jj}^{(C)}}\big)$
$=$ $-\frac{1}{2}(1-\frac{1}{c}){\rm E}_{q}({\tau^{-2}}){\sum_{j=1}^{p-1}}{P_{j}^{(C)}}{d_{jj}^{(C)}}-\frac{1}{2c}{\rm E}_{q}({\tau^{-2}}){\sum_{j=1}^{p-1}}{d_{jj}^{(C)}}$
$-\frac{p-1}{2}{\rm E}_{q}(\log{\tau^{2}})+\text{Const.},$
${\rm E}_{q}(\log p({\tau^{2}}|a))=$ $-\frac{1}{2}{\rm E}_{q}(\log a)-\frac{3}{2}{\rm E}_{q}(\log{\tau^{2}})-\frac{1}{2}{\rm E}_{q}({\tau^{-2}}){\rm E}_{q}(a^{-1})+\text{Const.},$
@@
$=$ ${\sum_{j=1}^{p-1}}[{P_{j}^{(C)}}E\left(\log{\pi_{j}}\right)+(1-{P_{j}^{(C)}}){\rm E}_{q}(\log(1-\pi_{j}))]+\text{Const.},$
${\rm E}_{q}(\log p(\pi))=$ $({\rho}_{1}-1){\sum_{j=1}^{p-1}}{\rm E}_{q}(\log\pi_{j})+({\rho}_{2}-1){\sum_{j=1}^{p-1}}{\rm E}_{q}(\log(1-\pi_{j}))+\text{Const.},$
@@
$=$ $-\frac{3}{2}{\rm E}_{q}(\log a)-A^{-1}{\rm E}_{q}(a^{-1})+\text{Const.},$
@@
$+(\alpha_{{\tau^{2}}}-1){\rm E}_{q}(\log{\tau^{2}})+\beta_{{\tau^{2}}}E({\tau^{-2}})+\text{Const.},$
@@
$-{\rm E}_{q}(\log q(a))=$ $-\log\Big(A^{-1}+{\rm E}_{q}({\tau^{-2}})\Big)+2{\rm E}_{q}(\log a)+\Big(A^{-1}+{\rm E}_{q}({\tau^{-2}})\Big){\rm E}_{q}(a^{-1})+\text{Const.}$
$=$ $\log\Big(A^{-1}+{\rm E}_{q}({\tau^{-2}})\Big)+\text{Const.}$
@@
$=$ $(\mathbf{y}-M_{\boldsymbol{\xi}})^{\top}\mathbf{X}\text{diag}({P^{(B)}}){\boldsymbol{\beta}}-\frac{1}{2}{\boldsymbol{\beta}}^{\top}\left[S_{X}^{\boldsymbol{\xi}}\odot\Omega+{\rm diag}({\rm E}_{q}(\alpha))\right]{\boldsymbol{\beta}}+\text{Const.},$
@@
$\Omega={\rm E}_{q}({\boldsymbol{\gamma}}{\boldsymbol{\gamma}}^{T})=({P^{(B)}})({P^{(B)}})^{\top}+{\rm diag}({P^{(B)}})({I_{p}}-{\rm diag}({P^{(B)}})).$
@@
$=$ $\gamma_{j}\Big[(y-M_{\boldsymbol{\xi}})^{T}X_{j}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(B)j}-\frac{1}{2}S_{X_{jj}}^{\xi}{{d_{jj}^{(B)}}}$
$-\frac{1}{2}\sum_{i\neq j}P_{i}^{(B)}S_{X_{ij}}^{\boldsymbol{\xi}}d_{ij}^{(B)}+{\rm E}_{q}\big(\log\big(\dfrac{\pi_{j}}{1-\pi_{j}}\big)\big)\big]+\text{Const.},$
@@
$=$ $-M_{\boldsymbol{\xi}}\mathbf{X}{P^{(B)}}{\boldsymbol{\mu}}_{{\boldsymbol{\beta}}(B)}-\frac{1}{2}\text{tr}[D_{{\boldsymbol{\beta}}}^{(B)}(S_{X}^{\boldsymbol{\xi}}\odot\Omega)]$
@@
$=$ $\frac{1}{2}{\sum_{j=1}^{p-1}}{\rm E}_{q}(\log\alpha_{j})-\frac{1}{2}\text{tr}\big[D_{{\boldsymbol{\beta}}}^{(B)}\text{diag}({\rm E}_{q}(\alpha))\big]\Big]+\text{Const.},$
@@
$=$ ${\sum_{j=1}^{p-1}}\big[(a_{j}-1){\rm E}_{q}(\log\alpha_{j})-b_{j}\frac{a_{j}+1/2}{b_{j}+\frac{d_{jj}}{2}}\big]+\text{Const.},$
@@
$=$ ${\sum_{j=1}^{p-1}}\left[{P_{j}^{(B)}}{\rm E}_{q}(\log\pi_{j})+(1-{P_{j}^{(B)}}){\rm E}_{q}(\log(1-\pi_{j}))\right]+\text{Const.},$
${\rm E}_{q}(\log p(\pi))=$ $({\rho}_{1}-1){\sum_{j=1}^{p-1}}{\rm E}_{q}(\log\pi_{j})+({\rho}_{2}-1){\sum_{j=1}^{p-1}}{\rm E}_{q}(\log(1-\pi_{j}))+\text{Const.},$
@@
$-{\rm E}_{q}(\log q(\boldsymbol{\alpha}))=$ $-{\sum_{j=1}^{p-1}}(a_{j}-1/2){\rm E}_{q}(\log\alpha_{j})-{\sum_{j=1}^{p-1}}(a_{j}+1/2)\log(b_{j}+d_{jj}/2)+\text{Const.}$
@@
Based on the results of the simulation study, we present a guideline for choosing a suitable prior in Table 4. The best model is suggested in each of the low- and high-dimensional cases, based on the sparsity, estimation, and prediction evaluation criteria.
For low-dimensional settings, the recommended approach is consistent across objectives: the Bernoulli-VB or Laplace-VB methods are suggested for enforcing sparsity, performing parameter estimation, and generating predictions. In contrast, for high-dimensional problems, the recommendation shifts to favor the CS-VB and Laplace-VB methods.
In general, the choice of prior depends critically on the dimensionality of the data. The Laplace-VB method stands as a good choice for all dimensions. The Bernoulli-VB emerges as a robust and simpler default for low-dimensional problems where explicit sparsity modeling is tractable. For high-dimensional inference, priors incorporating the CS mechanism become essential. Therefore, practitioners should first assess the dimension of their covariate space before applying the corresponding prior pairs from this guideline to ensure optimal model performance in sparsity recovery, estimation accuracy, and predictive quality.
@@
Sparsity Laplace & Bernoulli CS & Laplace
Estimation Bernoulli & Laplace Laplace & CS
Prediction Bernoulli & Laplace CS & Laplace